I need help with this question I was set for homework:
5/7(x-2) - 3/4(x+3) = -4
I've been trying to get rid of the denominator by finding the lowest common multiple, but I don't think that's correct.
Thanks in advance.

Answer by nerdybill(7384)

5/7(x-2) - 3/4(x+3) = -4
.
To get rid of the denominators, multiply both sides by 28:
20(x-2) - 21(x+3) = -112
20x-40 - 21x - 63 = -112
-40 - x - 63 = -112
-x - 103 = -112
-x = -9
x = 9